help please

Three red and three blue flags are arranged randomly along
a wire. What is the probability that the six flags alternate in
color?
A. 1/20
B. 1/10
C. 1/4
D. 1/2

Answer by sudhanshu_kmr(1152)

total no. of ways = 20
no. of ways when flags alternate is color = 2
probability = 2/20 = 1/10


ans: B. 1/10